Derek can deposit $211.00 per month for the next 10 years into an account at Bank A. The first deposit will be made next month. Bank A pays 12.00% and compounds interest monthly. Derek can deposit $2,450.00 per year for the next 10 years into an account at Bank B. The first deposit will be made next year. Bank B compounds interest annually. What rate must Bank B pay for Derek to have the same amount in both accounts after 10 years?

Please show how to answer step by step using a financial calculator
Answer #1

For Bank A, it is ordinary annuity with

A = $211

n = 12

t = 10

r = 12% compounded monthly

So, FV = A*((1+r/n)^(n*t) - 1)/(r/n) = 211*((1+0.12/12)^120 - 1)/(0.12/12) = $48538.16

For Bank B, it is also an ordinary annuity with

A = $2450

t = 10 year

FV = same as Bank A = $48538.16

So, FV = A*((1+r)^(t) - 1)/(r) => 48538.16 = 2450*((1+r)^10 - 1)/r

=> r = 14.50%

Bank B must pay a rate of 14.50%